# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ 4 Theoretical Analysis
In this section, we have given theoretical guarantees for the expected value of the regret and cumulative regret, but we have not mentioned high-probability bounds. Nevertheless, high-probability bounds can be easily derived from Markov’s inequality. In fact, for a non-negative random variable $X$, since Markov’s inequality gives $\mathbb{P}(X>a)\leq\mathbb{E}[X]/a$, if $a=\delta^{-1}\mathbb{E}[X]$, we can show that $X\leq\delta^{-1}\mathbb{E}[X]$ with probability at least $1-\delta$. As a result, for a non-negative value $J$ satisfying $\mathbb{E}[X]\leq J$, we can show that $X\leq\delta^{-1}J$ with probability at least $1-\delta$. Therefore, due to the non-negativity of the regret and cumulative regret, the following theorem holds by Theorem 4.1 and 4.2.
@@
Let $\delta\in(0,1)$. Under the assumption in Theorem 4.1, for any $t\geq 1$, the following inequality holds with probability at least $1-\delta$:
@@
where $h_{i}(\cdot)$ and each coefficient are given in Theorem 4.1. In addition, for the index $\hat{t}$ given by (4.3), the following holds with probability at least $1-\delta$:
@@
Although Theorem 4.4 gives high-probability bounds, these inequalities are not tight. In fact, the right-hand side of each inequality contains the inverse of $\delta$. On the other hand, most high-probability bounds based on GP-UCB such as Srinivas et al., (2010) use $\log(\delta^{-1})$ instead of $\delta^{-1}$. Therefore, making the bound of Theorem 4.4 tighter is one of our future work.
@@ 5 Numerical Experiments
In this section, we verify the performance of the proposed method using synthetic data functions and real data on the carrier lifetime values of silicon ingots used in solar cells. In all experiments, a GP model with the zero mean function was used as the surrogate model. Detailed experimental settings and experiments not described in the main body are described in Appendix D.
@@ 5.1 Synthetic Function
The input space $\mathcal{X}\times\Omega$ was a subset of $[-M,M]^{d}\times[-M,M]^{d}\equiv[-M,M]^{D}$ with each coordinate equally spaced into $s$ grid points. In this experiment, three different values of $(M,D,s)$ were considered: $(5,2,50)$, $(2.5,4,15)$, and $(2,6,7)$. As black-box functions, when $D=2$, $f$ was the sample path from a GP (2D synthetic function), and when $D=4$, $f(x_{1},x_{2},w_{1},w_{2})=f_{{\rm H}}(x_{1}+w_{1},x_{2}+0.5w_{2})$ was used (4D synthetic function), where $f_{{\rm H}}(a,b)$ is Himmelblau’s function with translation and scalar multiplication. When $D=6$, $f$ was the sum of sample paths $f_{1},\ldots,f_{4}$ from four independent GPs (6D synthetic function), where $f_{1}$, $f_{2}$, $f_{3}$ and $f_{4}$ depend only on $(x_{1},x_{2},x_{3})$, $(x_{2},x_{3},w_{1})$, $(x_{3},w_{1},w_{2})$ and $(w_{1},w_{2},w_{3})$, respectively. In each setting, the kernel function $k(\cdot,\cdot)$ of GPs, which is the surrogate model, was used as follows:
@@
Among these settings, only the 2D synthetic function setting correctly represented the true black-box function, and the remaining two settings were used in which the surrogate model did not match the black-box function. In addition, the experiment was conducted under the condition that the observations contained independent noise from the normal distribution with mean 0 and variance $10^{-6}$. As robustness measures, the following three cases were considered using the probability mass function $p({\bm{w}})$ of $w$ defined in each setting:
@@
Weighted sum of the expectation measure and mean absolute deviation,
@@
In the 2D, 4D and 6D synthetic function settings, $(h,\alpha)$ was set to $(0.5,1)$, $(0.18,4)$ and $(2,8)$, respectively. In this experiment, only the acquisition function was changed, and the evaluation index was set to regret $r_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$. For comparison, we compared nine methods including the proposed method: random sampling (Random), uncertainty sampling (US), Bayesian quadrature (BQ) (Beland and Nair,, 2017), BPT-UCB (Iwazaki et al., 2021a, ), BPT-UCB (fix), BBBMOBO (Inatsu et al., 2024a, ), BBBMOBO (fix), proposed method (Proposed), and Proposed (fix). Here, Random randomly selected $({\bm{x}}_{t},{\bm{w}}_{t})$, and US selected $({\bm{x}}_{t},{\bm{w}}_{t})$ by maximizing $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. For the other seven methods, ${\bm{x}}_{t}$ was selected by each acquisition function, and for the five methods other than BPT-UCB and BPT-UCB (fix), ${\bm{w}}_{t}$ was selected by (3.5). The method for selecting ${\bm{w}}_{t}$ in BPT-UCB and BPT-UCB (fix) is described in Appendix D. BQ and BPT-UCB are methods proposed for EXP and PTR, respectively. BBBMOBO is originally a method for Pareto optimization for multiple robustness measures, but it can also be applied to a single robustness measure. The trade-off parameters used in BPT-UCB, BBBMOBO, and Proposed were theoretical values. On the other hand, for methods marked with (fix), fixed smaller values than the theoretical values were used. Under these settings, we randomly selected one initial value and ran the algorithm until the number of iterations reached 300. This was repeated 100 times, and the average $r_{t}$ was calculated for each iteration. From Figure 2, it can be confirmed that Random and US, which are not acquisition functions for maximizing robustness measures, do not perform well in any setting. While BQ, BPT-UCB and BPT-UCB (fix) work effectively for EXP or PTR, the performance of EXP-MAE in the 2D synthetic function is not sufficient. This is because these are not acquisition functions for EXP-MAE. For BBBMOBO and Proposed, and BBBMOBO (fix) and proposed (fix), the results tend to be similar, because the only difference is whether ${\bm{x}}_{t}$ is set to $\tilde{\bm{x}}_{t}$ or (3.4) is used, except for the trade-off parameters. For BBBMOBO (fix) and proposed (fix), since the trade-off parameters are smaller than the theoretical values, practical performance tends to be high, and good performance is achieved in many settings. However, the regret has not been reduced completely in the EXP of the 4D synthetic function. One of the reasons for this is that the surrogate model cannot correctly express the true function, and furthermore, since the trade-off parameters are small, the search is not sufficient, as a result, resulting in being stuck in a local solution. In contrast, in Proposed, since the trade-off parameters are random, they can take large values, and as a result, there are opportunities to escape from the local solution, which suggests that performance is improved. Therefore, this is one of the advantages of the trade-off parameters, which are random, apart from the theoretical guarantee. The trade-off parameters of BBBMOBO are large on the order of $O(\log(t|\mathcal{X}\times\Omega|))$. Therefore, since they are often larger than the values of $\beta_{t}$ in Proposed, and are more conservative, Proposed performs better except for the 4D synthetic function (EXP). Overall, it can be confirmed that Proposed performs at least as well as the comparison methods in most settings.
@@
Figure 2: Average regrets for each method over 100 simulations. The top, middle and bottom rows show the results for 2D, 4D and 6D synthetic function settings, respectively, while the left, middle and right columns show the results for EXP, PTR and EXP-MAE, respectively. The length of the error bars represents two times the standard error.

### v2

@@ 4 Theoretical Analysis
Theorem 4.3 states that when considering the expectation measure as a robustness measure, an upper bound on $\mathbb{E}[r_{t}]$ is given, while for other robustness measures, estimation of $\hat{t}$ is still required. If the number of samples $M$ generated from the posterior distribution is sufficiently large, the estimated $\hat{t}$ can be expected to be close to the true $\hat{t}$, but whether it is possible to accurately derive the true $\hat{t}$ without estimation remains an important direction for future work. Nevertheless, in the synthetic function experiments conducted in Section 5.1, $\mathbb{E}[r_{t}]$ is used as the evaluation metric instead of $\mathbb{E}[r_{\hat{t}}]$, but the behavior of the convergence of $\mathbb{E}[r_{t}]$ to 0 in the proposed method has been confirmed in all settings. Therefore, similarly to how an upper bound for $\mathbb{E}[r_{t}]$ is given by Theorem 4.3, it is expected that some kind of upper bound holds for $\mathbb{E}[r_{t}]$ in other robustness measures.
In this section, we have presented theoretical guarantees for the expected value of both the regret and the cumulative regret. However, we have not addressed high-probability bounds. Such bounds can nonetheless be readily derived via Markov’s inequality. In fact, for a non-negative random variable $X$, Markov’s inequality yields $\mathbb{P}(X>a)\leq\mathbb{E}[X]/a$. By setting $a=\delta^{-1}\mathbb{E}[X]$, we obtain that $X\leq\delta^{-1}\mathbb{E}[X]$ with probability at least $1-\delta$. Hence, for any non-negative value $J$ satisfying $\mathbb{E}[X]\leq J$, we can conclude that $X\leq\delta^{-1}J$ with probability at least $1-\delta$. Applying this argument to the regret and cumulative regret, which are both non-negative, and invoking Theorems 4.1 and 4.2, we obtain the following result:
@@
Let $\delta\in(0,1)$. Under the assumptions of Theorem 4.1, for any $t\geq 1$, the following inequality holds with probability at least $1-\delta$:
@@
where $h_{i}(\cdot)$ and all coefficients are as defined in Theorem 4.1. In addition, for the index $\hat{t}$ defined in (4.4), the following inequality holds with probability at least $1-\delta$:
@@
Although Theorem 4.4 provides high-probability bounds, these bounds are not tight with respect to $\delta$. In fact, the right-hand sides of both inequalities depend on $\delta^{-1}$. In contrast, most high-probability bounds based on the GP-UCB framework—such as those in Srinivas et al., (2010)—involve a $\log(\delta^{-1})$ term instead of $\delta^{-1}$. Therefore, deriving tighter high-probability bounds with respect to $\delta$ than those in Theorem 4.4 remains an important direction for future work. Nevertheless, $\delta$ is a probability parameter specified in advance by the user and is a constant. Therefore, when considering the order with respect to $t$ by treating $\delta$ as a constant, the high-probability bound in the proposed method is tighter than that in GP-UCB-based methods such as Srinivas et al., (2010). For example, the standard BO on the input space $\mathcal{X}$ with no input uncertainty is equivalent to maximizing the expectation measure on $\mathcal{X}\times\Omega$ using $|\Omega|=1$. From Theorem 1 of Srinivas et al., (2010), the order of the high-probability upper bound for the cumulative regret is $\sqrt{t\gamma_{t}\log t}$, whereas in the proposed method, it is $\sqrt{t\gamma_{t}}$ according to Corollary 4.1. Furthermore, when considering the expectation measure as a robustness measure in the presence of input uncertainty, according to Theorem 1 and 2 in Inatsu et al., 2024a , at time $t$ the order of the high-probability bound for $r_{t^{\prime}}$ is $\sqrt{t^{-1}\gamma^{2}_{t}}$, where $t^{\prime}\leq t$. On the other hand, the high-probability bound for $r_{t}$ in our proposed method is $\sqrt{t^{-1}\gamma_{t}}$. However, as given in Table 1, the definitions of $r_{t}$ in Srinivas et al., (2010); Inatsu et al., 2024a and this study differ slightly, so it is important to note that these high-probability bounds cannot be directly compared.
4.1 Extension to Other Settings
In this section, we provide some results for uncontrollable settings. Furthermore, we provide an overview of methods and results for extending the proposed method to continuous spaces, and explain issues involved. Details on the extension to continuous spaces and to general uncontrollable settings are given in Appendix A and B.
Uncontrollable Setting when $\mathcal{X}$ and $\Omega$ are Finite
We consider the case of uncontrollable settings, i.e., ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow the distribution $P^{\ast}$ and cannot be controlled even during optimization. We assume that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ are mutually independent. The difference between the proposed algorithm in this setting and Algorithm 1 is only in how ${\bm{w}}_{t}$ is obtained. Specifically, in this setting, the acquisition function (3.5) cannot be used, and ${\bm{w}}$ sampled from $P^{\ast}$ becomes ${\bm{w}}_{t}$ at time $t$. We give the pseudo-code for the case where $\mathcal{X}$ and $\Omega$ are finite in the uncontrollable setting in Algorithm 2. When Algorithm 2 is performed, similar theorems as in the simulator setting hold.
Theorem 4.5.
Assume that the regularity assumption, Assumption 4.1 and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t}$, $\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\Omega=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(J)}\}$, $p_{j}=\mathbb{P}({\bm{w}}={\bm{w}}^{(j)})$ and $p_{min}=\min_{1\leq j\leq J}p_{j}$, and assume that $p_{min}>0$. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, under the uncontrollable setting, if Algorithm 2 is performed, the following holds:
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following inequality holds:
where $C$ is $8$ for the mean absolute deviation measure and $4$ for the other measures.
Theorem 4.6.
Under the assumptions of Theorem 4.5, the following holds:
where $\hat{t}$ is given by (4.4), and the function $h_{i}(\cdot)$ and all coefficients are as defined in Theorem 4.5. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
where $C$ is given in Theorem 4.5. Furthermore, in the expectation measure, the following holds:
Here, note that by using Theorems 4.5 and 4.6, and Markov’s inequality, high-probability bounds similar to those in Theorem 4.4 can also be obtained in this setting.
Extension to Continuous Spaces
In this section, we provide an overview of the extension of the proposed method to continuous spaces. In Algorithm 1 or 2, $\beta_{t}$ is defined using $|\mathcal{X}\times\Omega|$. Therefore, if $\mathcal{X}$ or $\Omega$ is a continuous space, this definition cannot be used directly. In this case, discretization of the continuous space is necessary, and to perform a theoretically valid discretization, additional assumptions on the differentiability of $f$, such as Assumption 2.1 in Takeno et al., (2023), are required. In fact, in Appendix A, the proposed method is extended to continuous spaces under similar assumptions. However, this approach has two issues. First, to propose a theoretically valid method, it is necessary to increase the number of discretized points according to time $t$, resulting in $\beta_{t}$ needing to be large according to time $t$, as in Theorem A.1. Thus, one of the advantages of the proposed method, namely that $\beta_{t}$ does not need to increase according to time $t$, is lost. Second, theoretically valid discretization requires constants such as $a_{1}$ and $b_{1}$ in Assumption A.1, and these are used to define $\beta_{t}$. However, although these constants $a_{1}$ and $b_{1}$ are determined by the kernel function, their actual values are difficult to calculate, resulting in the need for estimation or heuristic adjustment of these values. Therefore, the advantage of the proposed method, which eliminates the need for estimation or tuning of $\beta_{t}$, is also lost. Although an extension to continuous space is provided in Appendix A, the two issues mentioned above remain, and resolving these issues is an important direction for future work.
Algorithm 2 RRGP-UCB for robustness measures in the uncontrollable setting when $\mathcal{X}$ and $\Omega$ are finite.
GP prior $\mathcal{GP}(0,\ k)$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$
   Compute $Q_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (3.4)
   Generate $\bm{w}_{t}$ from $P^{\ast}$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
end for
@@ 5 Numerical Experiments
In this section, we verify the performance of the proposed method using both synthetic benchmark functions and real-world data on carrier lifetime values of silicon ingots used in solar cells. In all experiments, a GP model with a zero mean function is used as the surrogate model. Further details regarding the experimental settings, as well as additional experiments not included in the main text, are described in Appendix D.
@@ 5.1 Synthetic Function
The input space $\mathcal{X}\times\Omega$ was defined as a subset of $[-M,M]^{d}\times[-M,M]^{d}\equiv[-M,M]^{D}$; there each coordinate was uniformly discretized into $s$ grid points. Three different configurations of $(M,D,s)$ were considered in the experiments: $(5,2,50)$, $(2.5,4,15)$, and $(2,6,7)$. The black-box function $f$ varied depending on the dimension $D$. When $D=2$, $f$ was a sample path drawn from a GP, referred to as the 2D synthetic function. When $D=4$, the black-box function was defined as $f(x_{1},x_{2},w_{1},w_{2})=f_{{\rm H}}(x_{1}+w_{1},x_{2}+0.5w_{2})$ (4D synthetic function), where $f_{{\rm H}}(a,b)$ is Himmelblau’s function with translation and scaling. For the case $D=6$, $f$ was defined as the sum of four independent GP sample paths $f_{1},\ldots,f_{4}$ (6D synthetic function), where the dependencies of each function were specified as follows: $f_{1}$ on $(x_{1},x_{2},x_{3})$, $f_{2}$ on $(x_{2},x_{3},w_{1})$, $f_{3}$ on $(x_{3},w_{1},w_{2})$, and $f_{4}$ on $(w_{1},w_{2},w_{3})$. In each setting, the kernel function $k(\cdot,\cdot)$ used in the GP surrogate model was defined as follows:
@@
Among these settings, only the 2D synthetic function setting used a surrogate model that matched the true black-box function. The remaining two settings intentionally introduced a mismatch between the surrogate model and the black-box function. In addition, all experiments were conducted under the assumption that observations were corrupted by independent Gaussian noise with mean zero and variance $10^{-6}$. To evaluate robustness, three robustness measures were considered, using the probability mass function $p({\bm{w}})$ of ${\bm{w}}$ defined for each setting:
@@
Weighted sum of expectation measure and mean absolute deviation,
@@
In the 2D, 4D, and 6D synthetic function settings, the values of $(h,\alpha)$ were set to $(0.5,1)$, $(0.18,4)$, and $(2,8)$, respectively. In this experiment, only the acquisition function was changed, and the evaluation metric was the regret $r_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$. To compare performance, nine methods were evaluated, including the proposed method: random sampling (Random), uncertainty sampling (US), Bayesian quadrature (BQ) (Beland and Nair,, 2017), BPT-UCB (Iwazaki et al., 2021a, ), BPT-UCB (fixed), BBBMOBO (Inatsu et al., 2024a, ), BBBMOBO (fixed), the proposed method (Proposed), and Proposed (fixed). The random method selected $({\bm{x}}_{t},{\bm{w}}_{t})$ uniformly at random, while US selected $({\bm{x}}_{t},{\bm{w}}_{t})$ by maximizing $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. For the remaining seven methods, ${\bm{x}}_{t}$ was selected according to each method’s acquisition function. For all methods except BPT-UCB and BPT-UCB (fixed), ${\bm{w}}_{t}$ was selected using (3.5). The method for selecting ${\bm{w}}_{t}$ in BPT-UCB and BPT-UCB (fixed) is described in Appendix D. BQ and BPT-UCB were originally proposed for the EXP and PTR measures, respectively. Although BBBMOBO was designed for Pareto optimization over multiple robustness measures, it can be applied to a single robustness measure as well. The trade-off parameters used in BPT-UCB, BBBMOBO, and Proposed were set to theoretical values. In contrast, the methods marked with (fixed) used fixed values smaller than the theoretical ones. Note that in the EXP setting, BBBMOBO (fixed) and Proposed (fixed) are the same (see Appendix D for details). Under these settings, a single random initial point was selected, and the algorithms were run for 300 iterations. This process was repeated 100 times, and the average $r_{t}$ was calculated at each iteration. As shown in Figure 2, Random and US, which are not designed to maximize robustness measures, performed poorly in all settings. BQ, BPT-UCB, and BPT-UCB (fixed) were effective for EXP and PTR, but their performance for EXP-MAE in the 2D synthetic function was insufficient. This is because these methods are not tailored for EXP-MAE. For BBBMOBO and Proposed, as well as BBBMOBO (fixed) and Proposed (fixed), performance tended to be similar. This is because the only difference lies in whether ${\bm{x}}_{t}$ is set to $\tilde{\bm{x}}_{t}$ or selected via (3.4), aside from the trade-off parameters. For BBBMOBO (fixed) and Proposed (fixed), using smaller-than-theoretical trade-off parameters led to improved practical performance, achieving favorable results in many settings. However, in the 4D synthetic function under the EXP measure, regret was not fully reduced. One reason is that the surrogate model fails to correctly express the true black-box function. Furthermore, small trade-off parameters limit exploration, often resulting in convergence to local optima. In fact, Figure 5 of Srinivas et al., (2010) shows the results of BO (Mean Only) using only the posterior mean, and Figure 5 of Inatsu et al., 2024a compares the differences in $\beta$ for the expectation measure, suggesting that in both cases, small $\beta$ tends to lead to local optima. In contrast, the Proposed method, by employing random trade-off parameters, occasionally explores more broadly. This increases the likelihood of escaping local solutions and, consequently, improves performance. This demonstrates a key advantage of using randomly varying trade-off parameters beyond the theoretical guarantees. The trade-off parameters in BBBMOBO are on the order of $O(\log(t|\mathcal{X}\times\Omega|))$. Since they are often larger and more conservative than the $\beta_{t}$ values used in Proposed, Proposed generally outperformed BBBMOBO, except in the 4D synthetic function (EXP) setting. Overall, the Proposed method performed comparably to or better than the baseline methods across most settings.
@@
Figure 2: Average regret across 100 simulations for each method. Top, middle, and bottom rows correspond to 2D, 4D, and 6D synthetic function settings, respectively. Left, center, and right columns show results for EXP, PTR, and EXP-MAE, respectively. Since BBBMOBO (fixed) and Proposed (fixed) are the same in EXP, only Proposed (fixed) is displayed in the figures on the left column. Error bars represent twice the standard error.
5.2 Synthetic Function Experiments under Uncontrollable Settings
In this section, we changed the experiment in Section 5.1 to the uncontrollable setting, where ${\bm{w}}_{t}$ cannot be selected and is obtained randomly according to the probability mass function $p({\bm{w}})$. The method for selecting ${\bm{x}}_{t}$ was the same as in Section 5.1. As in Section 5.1, Figure 3 shows that the proposed method outperforms the comparison methods in most settings. On the other hand, in the experiments in Section 5.1, the regret of Proposed (fixed) did not decrease in the 4D synthetic function (EXP), while in Figure 3, the regret decreased more than Figure 2. This is because ${\bm{w}}$ could not be selected, and as a result of random sampling, more exploration was performed, making it possible to avoid local solutions.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Average regrets across 100 simulations under the uncontrollable setting for each method. Top, middle, and bottom rows correspond to 2D, 4D, and 6D synthetic function settings, respectively. Left, center, and right columns show the results for EXP, PTR, and EXP-MAE, respectively. Since BBBMOBO (fixed) and Proposed (fixed) are the same in EXP, only Proposed (fixed) is displayed in the figures on the left column. Error bars represent twice the standard error.